OpenStudy (anonymous):

ok im not sure how to write this any other way that makes it easier to understand. simplify (3x^29y^23/96x^4y^13)^(-1/5)

OpenStudy (anonymous):

start inside the bracklets (32x^4y^13)/(x^29y^23) ^(1/5) then long divide by 5 remainders stay under the fifth root

OpenStudy (anonymous):

\[\frac{2}{x^{5}y^{2}}\]

OpenStudy (anonymous):

do you take 32^(-1/5)

OpenStudy (anonymous):

i'll write out the steps if you would like

OpenStudy (anonymous):

\[\sqrt[5]{(\frac{32x^{4}y^{13}}{x^{29}y^{23}})} =\sqrt[5]{(\frac{32}{x^{25}y^{10}})}=\frac{2}{x^{5}y^{2}}\]